OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (danjs):

multiply the first equation by 3, then add them together to eliminate y

OpenStudy (danjs):

3x - 3y = -9 (3 times the first equation) x + 3y = 5 ------------- add them together 4x = -4 no more y variable like that

OpenStudy (danjs):

divide both sides of the remaining equation by 4 x = ...

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

is that the answer?